OpenStudy (anonymous):

If g(x) = x2 − 1, find g(3).

OpenStudy (anonymous):

it is 8

OpenStudy (anonymous):

plug in 3 for x g(3) = 3^2 - 1 g(3) = 9 - 1 g(3) = 8

OpenStudy (anonymous):

thanks